Please help me solve this equation: Find the volume of a pyramid with height 3 cm and rectangular base with dimensions 1cm and 2cm.
Answer by mananth(16946)

Volume V of a pyramid = 1/2* A * h
Arae = 2*1 = 2
h= 3
V= 1/2 * 2 * 3
V= 3 cu. cm